The ages of three brothers combined is 83 years. The oldest brother is twice the age of the middle and the middle brother is 5 years older than the youngest.
Answer by mananth(16948)

The ages of three brothers combined is 83 years. The oldest brother is twice the age of the middle and the middle brother is 5 years older than the youngest.
youngest ----------x
middle -------------(x+5)
eldest-------------2(x+5)
x+(x+5)+2(x+5)=83
x+x+5+2x+10=83
4x+15=83
4x=68
/4
x=17
17,22, 44 are the ages